OpenStudy (whoopsmehnoknow):

I think i know what it is but i am not quite positive can someone help me plz. :> Two lines, A and B, are represented by equations given below: Line A: y = x − 4 Line B: y = 3x + 4 Which of the following shows the solution to the system of equations and explains why? (−3, −5), because the point satisfies one of the equations (−3, −5), because the point lies between the two axes (−4, −8), because the point satisfies both equations (−4, −8), because the point does not lie on any axis

OpenStudy (whoopsmehnoknow):

I believe it is B

OpenStudy (whoopsmehnoknow):

hello?

OpenStudy (photon336):

@WhoopsMehNoKnow do you know how to do this?

OpenStudy (whoopsmehnoknow):

no i do not my lesson is confusing.

OpenStudy (photon336):

so it's asking for the solution to the two equations

OpenStudy (photon336):

this means it's the point where the two lines intersect

OpenStudy (photon336):

what are your thoughts on this?

OpenStudy (whoopsmehnoknow):

mhm like one solution or infinite

OpenStudy (photon336):

one solution

OpenStudy (whoopsmehnoknow):

or two points on a graph

OpenStudy (whoopsmehnoknow):

ya

OpenStudy (photon336):

alright

OpenStudy (photon336):

y = x-4 and y = 3x+4

OpenStudy (photon336):

so what we do is we set these two equal to eachother x-4 = 3x+4

OpenStudy (whoopsmehnoknow):

mhm

OpenStudy (photon336):

now solve as we would any normal equation. show me how you do this

OpenStudy (whoopsmehnoknow):

k one sec i like to right things down as well

OpenStudy (photon336):

get back to me when you have an answer

OpenStudy (whoopsmehnoknow):

would i add 4 to both sides? first?

OpenStudy (photon336):

yep

OpenStudy (whoopsmehnoknow):

so it would be x - 4 + 4 = 3x + 4 + 4

OpenStudy (photon336):

good

OpenStudy (photon336):

keep going

OpenStudy (whoopsmehnoknow):

and 4 +4 cancel out on both sides?

OpenStudy (photon336):

-4+4 cancel because they are opposite signs

OpenStudy (whoopsmehnoknow):

ahh

OpenStudy (photon336):

x = 3x+8

OpenStudy (photon336):

now what do we do?

OpenStudy (whoopsmehnoknow):

umm add it out like a normal problem x = 3x + 8 ---> x =11x

OpenStudy (photon336):

you can only add or subtract like terms

OpenStudy (whoopsmehnoknow):

ohh

OpenStudy (photon336):

you would subtract three from both sides like this x-3x = 3x-3x+ 8 -2x = 8

OpenStudy (whoopsmehnoknow):

AHH

OpenStudy (photon336):

now how would we find x?

OpenStudy (whoopsmehnoknow):

one sec brain just went out

OpenStudy (photon336):

-2x = 8

OpenStudy (whoopsmehnoknow):

wait so then you do divide -2x into 8 as well

OpenStudy (photon336):

yeah we divide because we always do the opposite operation

OpenStudy (photon336):

-2x means -2 times x

OpenStudy (photon336):

so what is x?

OpenStudy (whoopsmehnoknow):

which =s -4x

OpenStudy (photon336):

yeah x = -4

OpenStudy (photon336):

now we need to find something else what is it?

OpenStudy (whoopsmehnoknow):

is it Y? to make the full solution?

OpenStudy (photon336):

exactly we need to find y

OpenStudy (photon336):

to do this it's easy just plug your value for x back into one of the original equations

OpenStudy (photon336):

now if your x is correct and it is an intersection point you should get the same y value for both equations, why do you think this has to be always true?

OpenStudy (whoopsmehnoknow):

to make it one solution? on a graph?

OpenStudy (photon336):

 say if we have to lines drawn like this

OpenStudy (photon336):

the lines intersect at how many points?

OpenStudy (whoopsmehnoknow):

two say like (0,0)

OpenStudy (photon336):

alright well the idea here is that the two lines intersect at 1 point. and at that point both lines have the same x and y values

OpenStudy (whoopsmehnoknow):

ohh

OpenStudy (photon336):

yeah so whatever x value you get when you plug it back into one equation you must get the same y value for both of them

OpenStudy (whoopsmehnoknow):

so we would do so for y=3x + 4 (-4x for x)?

OpenStudy (photon336):

yes

OpenStudy (whoopsmehnoknow):

so y = 3 (-4) + 4

OpenStudy (photon336):

yes continue

OpenStudy (whoopsmehnoknow):

y = -12 + 4? because you always do the parentheses first

OpenStudy (photon336):

exactly keep going

OpenStudy (whoopsmehnoknow):

thats were i get stuck on...i keep getting confused on what to do from there

OpenStudy (photon336):

so in the ( ) there's multiplication. you do that first

OpenStudy (photon336):

3(-4) = -12 then you get -12+8

OpenStudy (photon336):

which equals what?

OpenStudy (whoopsmehnoknow):

-4?

OpenStudy (whoopsmehnoknow):

-12 + 8 = -4? or is that wrong

OpenStudy (photon336):

oh sorry lol I made a mistake

OpenStudy (photon336):

it's -12+4

OpenStudy (whoopsmehnoknow):

ahhh then its -8

OpenStudy (whoopsmehnoknow):

y = -8

OpenStudy (photon336):

okay yes but

OpenStudy (photon336):

to drill the concept into your head

OpenStudy (photon336):

plug x = -4 into y = x-4

OpenStudy (whoopsmehnoknow):

mk

OpenStudy (whoopsmehnoknow):

so it would be y = -4x - 4 or just normal without the x

OpenStudy (photon336):

so wherever you see x put a -4 there

OpenStudy (whoopsmehnoknow):

ohh

OpenStudy (whoopsmehnoknow):

it would be y = -8?

OpenStudy (photon336):

yeah and the last thing. what do you notice about those answers?

OpenStudy (whoopsmehnoknow):

they both equal each other. y=-8 y=-8

OpenStudy (photon336):

that's the point

OpenStudy (whoopsmehnoknow):

so its like checking your work?

OpenStudy (photon336):

that's how you know that they intersect

OpenStudy (whoopsmehnoknow):

ahh makes since

OpenStudy (photon336):

yeah that too but it's to prove a point that for two lines to intersect they intersect at a point and must both have the same x and y values at that point.

OpenStudy (whoopsmehnoknow):

ok much much more easy my teacher was not explaining that to me

OpenStudy (whoopsmehnoknow):

so the answer is c

OpenStudy (photon336):

yeah that's the point of all this

OpenStudy (photon336):

exactly

OpenStudy (whoopsmehnoknow):

YAY thankyou soooo much i have been stuck on that one question for 3 hours (no joke)

OpenStudy (photon336):

no problem